1) A garage door opener is rated at 373 W. In 14.0 s of operation the opener can do ___x103 J of work.

 

2) An automobile engine exerts a force of 595 N to propel the car at 90.0 km/h on a level road. The power output of the engine at that time is ___x10W.


Use the information below to answer the following


A traditional, double-seat chairlift can move1200 people in 1.00 hrs (with an average mass of 60.0 kg) up a ski hill. A high-speed quad chair can move 2400 people in 30.0 minutes up the hill. The upper terminal (top of the hill) is 100 m higher than the lower terminal (bottom of the hill).

4) The amount of power the double-seat chair requires just to overcome gravity is ___x10W. 





5) An electric motor has a power rating of 1.60 kW and operates at 75% efficiency. The amount of work that it can do in 1.0 hrs is ___x10J.





1
Expert's answer
2021-07-19T09:49:28-0400

1)

"A=Pt=5.2\cdot 10^3~J,"

2)

"P=Fv=1.5\cdot 10^4~W,"

4)

"P=\frac{Fh}{2t}=\frac{1200mgh}{2t}=1\cdot 10^4~W,"

5)

"P=\eta P_0,"

"\frac At=\eta P_0,"

"A=\eta P_0t=4.32\cdot 10^6~J."